OpenStudy (sambam):

Please Solve for x: 11x^2-7=2-5x^2 Also if you can please show your work so I can learn how to to do the problem.

OpenStudy (anonymous):

okay so we can add 'like' terms... meaning 11x^2 and -5x^2 are 'like' terms

OpenStudy (uri):

Sign is going to change when it shifts there.

OpenStudy (anonymous):

so we add 5x^2 to both sides of the equation and we get 16x^2 - 7 = 2 then add 7 to both sides to get 16x^2 = 9 now we can take the square root of both sides to get [(4x)^2 = 16 btw...] 4x = 3 divide by 4 on both sides to get x = 3/4

OpenStudy (anonymous):

So in the end we have (5x - 6)(x - 1)

OpenStudy (anonymous):

@tyrell84 what?

OpenStudy (anonymous):

f 5x^2 - 11x + 6 = 5x^2 - 5x - 6x + 6 = (5x^2 - 5x) - (6x - 6) = 5x(x - 1) - 6(x - 1) = (x - 1)(5x - 6) anwser (5x - 6)(x - 1)

OpenStudy (anonymous):

it's 11x^2

OpenStudy (anonymous):

robz

OpenStudy (anonymous):

ur not doing the caulatur not rite

OpenStudy (anonymous):

it is \[11x^{2}\]not 11x

OpenStudy (anonymous):

bro i not going to ARUGE WIT HU

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

if you're going to troll, i would recommend xbox live, not openstudy

OpenStudy (anonymous):

nerd lol get out ur mom basement then u can try to get on sumone